OpenStudy (anonymous):

solve by elmination. x+8y=8 -6x+9y=9

sam (.sam.):

x+8y=8 times -6 -6x+9y=9 ---------------------- -6x-48y=-48 -6x+9y=9 -57y=-57 y=1 x+8y=8 x+8=8 x=0

OpenStudy (anonymous):

thanks

sam (.sam.):

yw

OpenStudy (anonymous):

You actually multiply it by 6 not -6